### v1

Abstract
Many studies have examined social determinants of health (SDoH) factors independently, overlooking their interconnected and intersectional nature. Our study takes a multifactorial approach to construct a neighborhood-level measure of SDoH and explores how neighborhood residency impacts care received by endometrial cancer patients in Massachusetts. We used a Bayesian multivariate Bernoulli mixture model to create and characterize neighborhood SDoH (NSDoH) profiles using the 2015–2019 American Community Survey at the census tract level (n=1478), incorporating 18 variables across four domains: housing conditions and resources, economic security, educational attainment, and social and community context. We linked these profiles to Massachusetts Cancer Registry data to estimate the odds of receiving optimal care for endometrial cancer using Bayesian multivariate logistic regression. The model identified eight NSDoH profiles. Profiles 1 and 2 accounted for 27% and 25% of census tracts, respectively. Profile 1 featured neighborhoods with high homeownership, above-median incomes, and high education, while Profile 2 showed higher probabilities of limited English proficiency, renters, lower education, and working-class jobs. After adjusting for sociodemographic and clinical characteristics, we found no statistically significant association between NSDoH profiles and receipt of optimal care. However, compared to patients in NSDoH Profile 1, those in Profile 2 had lower odds of receiving optimal care [OR = 0.77, 95% CI (0.56, 1.07)]. Our results demonstrate the interconnected and multidimensional nature of NSDoH, underscoring the importance of modeling them accordingly. This study also highlights the need for targeted interventions at the neighborhood level to address underlying drivers of health disparities, ensure equitable healthcare delivery, and foster better outcomes for all patients.
keywords:
Bayesian clustering; Endometrial Cancer; National Comprehensive Cancer Network; Social Determinants of Health;disparities; neighborhoods
1 Background
Social determinants of health (SDoH) are increasingly recognized as critical factors that shape healthcare outcomes by influencing access to care, treatment decisions, and overall patient experiences[1, 2, 3]. These determinants, including socioeconomic status, experiences of racism or discrimination, and neighborhood conditions, can introduce bias in the care that individuals receive, potentially leading to disparities in treatment and outcomes[4]. Endometrial cancer (EC), one of the most commonly occurring female cancers, serves as an example of how SDoH can influence care[5, 6]. Despite the existence of evidence-based treatment guidelines, disparities across the EC care continuum persist, particularly among ethnically and socioeconomically disadvantaged groups[7, 8, 9, 10, 11, 6, 5, 12, 13]. A comprehensive analysis of SDoH that transcends race and ethnicity would provide researchers and policymakers with a deeper understanding of how the social environment can influence a patient’s ability to engage with the healthcare system [11, 12].
In settings where individual-level SDoH measures are not readily available, researchers often rely on non-standardized composite scores of proxy variables. These variables correspond to different domains of SDoH, such as income, education, and employment[14, 15]. At the geographic level, SDoH has been operationalized using area-based measures derived from publicly available data sources such as the United States (US) Census and the American Community Survey (ACS). Two commonly used measures are the Yost Index and the Area Deprivation Index (ADI)[16, 17, 18]. The Yost Index, employed in national cancer surveillance, currently operates under the assumption that SDoH is one-dimensional. It inputs seven area-level variables (median household income, education, proportion of households below the 150% poverty line, median house value, median gross rent, unemployment rate, and working-class occupation). These variables are summarized at the census tract or block level through a Principal Component Analysis (PCA). Here, the primary component, which accounts for the most variance among the variables, is selected as the socioeconomic status (SES) index.[16, 19, 20]. The component loadings are used to calculate a composite SES score for each areal unit, and these are further ranked based on their SES scores and divided into quintiles, with quintile 1 representing the lowest SES and quintile 5 the highest.
The ADI, originally developed in 2002 to study general mortality disparities and hospitalizations in the US, expands the set of census-derived SDoH variables to 17. Previously, it operated under the same one-dimension assumption as Yost, but now incorporates a three-factor model to accommodate multi-dimensionality [21, 20, 18, 17]. Highly correlated variables are first clustered together and then defined along the following three domains: financial strength, economic hardship and inequality, and educational attainment [18]. As in the Yost Index, areal units are indexed based on the loadings of each of these three-factor domains, where higher percentile ranks indicate greater socioeconomic disadvantage. Boscoe et al. compared these two indices and found that despite being derived from the same data and heavily influenced by financial domains, they differ significantly in terms of transparency for reproducibility, distribution, and sensitivity to missing data[20]. Methodologically, these approaches share similarities in their use of principal components or factor analysis to reduce the number of highly correlated SDoH variables, but the domains derived are treated independently from one another and fall short of acknowledging the ways in which they intersect.
Recently, Kolak et al. introduced a new approach that further expands the list of census-derived SDoH variables for PCA, including variables like limited English proficiency [3]. These variables were reduced to four domains: socioeconomic advantage index, limited mobility index, urban core opportunity index, and mixed immigrant cohesion and accessibility index. Additionally, they clustered census tracts by characterizing them into profiles based on similar scoring across the four SDoH components using K-means clustering. While this approach does allow for clustering based on shared responses to the different domains, the clustering is based on a single summary score. Overlooking how each component-specific score contributes to the summary score ignores the heterogeneous ways in which area deprivation can manifest. Two neighborhoods could potentially be clustered together but experience different patterns of resource allocation, which could ultimately influence how those residents interact and are able to access healthcare services. New methodological approaches are needed to accommodate the interconnectedness and multifactorial structure of SDoH, inclusive of factors pertinent to vulnerable populations that are at greater risk of health disparities[20, 3, 14, 22].
Model-based clustering methods, specifically finite mixture models (FMM), are able to identify underlying patterns in heterogeneous populations from a wide set of interrelated variables. Lekkas et al. recently highlighted the methodological advantages of FMMs in neighborhood health research through the common application of Latent Class Analysis (LCA) and Latent Profile Analysis (LPA). These models are able to characterize neighborhood profiles based on a set of multivariate categorical or continuous data and examine their association with various health indicators [23]. Although LCA or LPA are well-suited to handle the multi-faceted nature of neighborhoods, typically, the number of latent patterns/profiles is unknown, and multiple model fitting and testing is required to determine the optimal number of classes. However, under a Bayesian framework, we are able to rely on data-driven algorithms to identify an interpretable and reproducible set of neighborhood SDoH (NSDoH) profiles with the added flexibility of incorporating prior information into parameter estimation [24]. These models have found great utility in nutrition, environmental health, and population genomics [25, 26, 27, 28, 29, 30, 31]. Yet, none have been applied to area-level exposures of SDoH.
In this study, we aim to generate and characterize new NSDoH profiles to better understand their impact on EC care in Massachusetts. Given binarized aggregate proportions from the ACS, we will implement a Bayesian multivariate Bernoulli mixture model(MBMM) to cluster neighborhoods that share similar SDoH characteristics. Neighborhoods are established based on census tract identifiers, as defined by the U.S. Census Bureau from 2015 to 2019. Once identified, we will use these profiles with patient-level data from the Massachusetts Cancer Registry to determine if the receipt of optimal care for EC differs by a patient’s residential neighborhood.
We organize the remaining sections of this paper as follows. In section 2, we describe our methodological approach. In section 3, we share our results of the MBMM and examine its association with EC treatment. Finally, in sections 4 and 5, we provide a discussion and concluding remarks.
2 Methods
2.1 Data Source
We used census tract (neighborhood) level aggregate data from the 2015-2019 five-year American Community Survey (ACS) estimates for the state of Massachusetts (MA). The data contained a total of 1478 census tracts that included approximately 6.8 million persons. Based on previous research, we selected eighteen(18) variables as indicators of NSDoH [20, 3]. Given that the selected NSDoH variables were aggregate proportions and the Bernoulli distribution models binary data, we dichotomized each variable based on whether that feature fell above $(x_{ij}=1)$ or below $(x_{ij}=0)$ the state median value. Table 1 details the median threshold determined for each of the SDoH features. Neighborhoods above that specific threshold were classified as high exposure. Neighborhoods below that specific threshold were classified as low exposure. Data was extracted from respective ACS tables via the tidycensus package implemented in the R software environment [32].
[…]
2.2 Multivariate Bernoulli Mixture Model (MBMM)
Neighborhood social determinants of health (NSDoH) profiles were identified using a fully Bayesian estimation of a multivariate Bernoulli mixture model (MBMM). This model clusters a population based on shared responses to a set of binary exposures. Let $i\in\{1,...,n\}$ index an individual census tract, where $n$ is the total number of census tracts or neighborhoods in Massachusetts. Let $\mathbf{x_{i}}=\{x_{i,1},..,x_{i,p}\}$ denote a vector of observed binary indicators, where $x_{i,j}=1$ denotes a high exposure (above state median threshold) to social determinant $j\in\{1,..,p\}$ in census tract $i$. Let $K$ denote the number of NSDoH profiles in the population. In practice, this number is typically not known a priori. To determine an appropriate number of profiles in the model, we overfit the model with $K_{max}$ clusters, which over-exceeds the true number of clusters and allows the observed data to drive the number of nonempty clusters estimated at the end of the Bayesian sampling algorithm [44]. The observed likelihood for the set of binary data $\mathbf{X}=\{0,1\}^{n\times p}$ under the MBMM is given by
$\mathcal{L}(\boldsymbol{\pi},\boldsymbol{\theta}|\mathbf{X})=\prod_{i=1}^{n}\biggl{\{}\sum_{k=1}^{K}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}\biggl{\}},$ (1)
where $\boldsymbol{\pi}=\{\pi_{1},...,\pi_{K}\}$ is the probability vector for membership of a given neighborhood into one of the $K$ NSDoH profiles. The probability matrix $\boldsymbol{\theta}=\{\theta_{1|1},...,\theta_{j|k}\}^{p\times K}$ summarizes the set of individual SDoH variables, where $\theta_{j|k}$ is the probability of a high exposure to variable $j$ given the neighborhood’s assignment to NSDoH profile $k$.
For estimation of these parameters, we augment the data by introducing a latent allocation variable $\boldsymbol{z_{i}}$, such that $P(z_{i}=k)=\pi_{k}$. Therefore, we consider the complete data $\{x_{i},z_{i}\}$ likelihood of the MBMM for computation:
$\mathcal{L}^{c}(\boldsymbol{\pi},\boldsymbol{\theta}|\mathbf{X},\mathbf{Z})=\prod_{i=1}^{n}\prod_{k=1}^{K}\biggl{\{}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}\biggl{\}}^{\mathbb{I}(z_{i}=k)}$ (2)
2.2.1 Posterior Computation
Model parameters are estimated via a Metropolis-coupled Markov chain Monte Carlo ($MC^{3}$) algorithm. An allocation sampler is used to determine the most probable number of NSDoH profiles. This is based on the maximum a posteriori number of nonempty clusters from the posterior distribution,$K_{map}$. Once $K_{map}$ is determined, we calculate the posterior mean of each of the model parameters. These posterior estimates are then used to describe each NSDoH profile $k$, based on the posterior distribution of $\boldsymbol{\theta_{\cdot|k}}=(\theta_{1|k},\ldots,\theta_{p|k})$, and assigning each MA census tract to the profile with the greatest posterior probability of membership. Further details on prior selection, model fit testing, and other cluster postprocessing procedures are provided in Supplemental Materials A.
2.2.2 Regression Analysis
Patient-level data was obtained from the Massachusetts Cancer Registry (MCR) of the Massachusetts Department of Public Health (MDPH). A total of 2412 records were collected for females, aged 18 years or older and diagnosed with endometrial cancer between 2015 and 2017. Cases were identified using the International Classification of Diseases for Oncology, third edition (ICD-O3) primary site (C54.1) and morphology codes for endometrial carcinoma [45]. The main outcome variable was adherence to National Comprehensive Cancer Network (NCCN) guidelines for the first course of treatment while accounting for guideline changes during the study period. These guidelines are based on tumor, grade, and stage, incorporating a combination of surgery, chemotherapy, and radiation as necessary. Adherence to NCCN treatment guidelines was treated as a dichotomous variable where we defined optimal care as the patient receiving therapy in accordance with NCCN guidelines. The main exposure was NSDoH profiles derived from the MBMM in 2.2 and assigned to each patient’s residential census tract at EC diagnosis. Associations were examined by adjusted odds ratio using a Bayesian logistic regression, adjusting for patient-level clinical and sociodemographic characteristics at cancer diagnosis (year of diagnosis, age, insurance status, and initial point of care facility type). Additionally, we conducted a supplementary analysis to examine the association between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). More details about the guidelines and regression analyses are provided in the Supplemental Materials Section B.
All codes for the MBMM, data analysis, and data wrangling, including ACS datasets, are available on GitHub at https://github.com/cbrodriguez01/ecbayesbinmix. All statistical analyses were conducted using the R software environment (version 4.3.1), and R packages (not exhaustive) tidycensus, ggplot2, tidyverse, table1, BayesBinMix, coda, readxl, stringr, brms, shiny, leaflet, tigris, and sf. The MBMM computations were run on the Harvard University Faculty of Arts and Sciences Research Computing (FASRC) Cannon cluster.
3 Results
3.1 NSDoH Profiles Results
Table 1 shows the NSDoH variables that were selected from the 2015-2019 ACS survey and clustered to characterize census tracts in Massachusetts. MA census tracts were mostly composed of owner-occupied households and high proportions of $\geq$ HS and $\geq$ Bachelor’s degree educational attainment. On average, across all census tracts, about 13% of households received SNAP benefits in the past 12 months, and 52% of the population aged 16 or older had non-white collar occupations.
The model identified eight NSDoH profiles. Figure 1 illustrates the estimated posterior probability of high exposure to each of the 18 NSDoH variables grouped by domain, given assignment to that specific NSDoH profile. Profile 1 represented 27% of MA census tracts. This profile is characterized by neighborhoods that are more likely to have owner-occupied housing (99.7%), income above the state median (95.8%), and higher than high school education (95.6%). It also had a lower probability of exposure to working-class employment (4.7%), participation in federal assistance programs (6.6%), and most social and community context indicators. Profile 2 represented the second largest cluster, containing 25% MA census tracts. This cluster had the highest probability of exposure to households with limited English proficiency (96.4%), female head of households (90.8%), and household crowding (84.1%). High probabilities of exposure were also identified for neighborhoods with a lack of complete plumbing (39.5%), renter-occupied housing (99.7%), no vehicle access (94.6%), working class (98.1%), and all three ethnic minority groups, especially the proportion of non-Hispanic Black (81%) and Hispanic/Latino residents (94.5%). This profile also had the lowest education levels, with almost all reporting less than HS education (98.4%), compared to all other NSDoH profiles. Overall, this profile exhibited characteristics of disadvantage and deprivation across all thematic domains. Profile 3 favored neighborhoods with high probabilities of exposure to housing conditions and resources variables (renter-occupied, no vehicle, household crowding), high proportions of ethnic minorities (51-83%), and educational attainment (HS and Bachelor’s degree education). Profile 4 shared similarities with NSDoH profile 1 along the housing conditions and resources domains but differed in economic security (more working class) and education (less Bachelor’s degree education). It also had the lowest probabilities of exposure for social and community context indicators compared to all other profiles. Profile 5 favored neighborhoods with high probabilities of exposure to owner-occupied housing, less than HS education, and working class, moderate probability of ethnic minorities (48-68%), and low probabilities for lack of complete plumbing (21.4%) and no vehicle (30.2%). Profile 6 favored neighborhoods with high probabilities of exposure to higher proportions of ethnic minority groups (85-95%), residents with limited English proficiency (94.5%), renter-occupied housing (98.9%), no vehicle in the household (96.5%), and SNAP benefits (90.8%). Although similar to Profile 2, a notable observation is a divide in educational attainment, particularly between those with less than HS education (85.8%) and those with at least a Bachelor’s degree education (96%). Profile 7 shared similarities with NSDoH profile 5, but differed in education and social and community context variables. Profile 7 had higher probabilities of exposure to those with high school and bachelor’s degree education and lower probabilities of exposure to the proportion of ethnic minority residents (29%-33%) and limited English proficiency (17%). Lastly, Profile 8 was characterized by neighborhoods with a higher probability of exposure to residents with less than HS education (47.1%), households living below the poverty line (84.4%), participation in federal assistance programs (82.7%), renter-occupied housing (98.7%), and working-class employment (80.8%), and a lower probability of exposure to households with income above the state median (10.2%). Neighborhoods in this profile had a 60% probability of being exposed to residents with a female head of household and close to 50% estimated probability for an above state median proportion of Hispanic/Latino or non-Hispanic Black residents.
3.2 Spatial Mapping of NSDoH
The MBMM profile assignments for each census tract were recorded and linked with census tract geographic data. We developed a web-based interactive map using Shiny to display the MBMM results, available at https://mhn38j-carmen-rodriguez.shinyapps.io/nsdoh_profiles_app_urban/. In this map, we also added a layer delineating 2020 Urban Boundaries obtained from the Massachusetts Bureau of Geographic Information Data Hub [https://gis.data.mass.gov] to further describe the NSDoH profiles.
Neighborhoods that were assigned to Profile 1 were found predominantly in surrounding areas outside of Boston (e.g. Middlesex and Essex counties). Examining the overlaid urban boundaries reveals that most census tracts in this profile are located within urbanized areas. Additionally, the average proportion of Non-Hispanic White individuals in these census tracts was 85.9%, which further corroborates the results of our model for social and community context indicators. Neighborhoods assigned to Profile 2 were identified in areas with large immigrant populations (e.g., Lawrence, Lynn, Chelsea, Springfield) and a high average proportion of Hispanic/Latino residents. Neighborhoods in Profile 3 were primarily in areas with a known university presence and an average median age of 35 years old (e.g., Cambridge, Amherst, Brookline). Neighborhoods in Profile 4 were primarily in areas south and west of Boston (e.g. Berkshires, Plymouth). Neighborhoods in Profile 5 (e.g. Saugus, Winchendon, Brockton) were close in proximity to neighborhoods in Profile 2, and they shared similar age and race and ethnicity distributions in their population. Neighborhoods in profile 6 were primarily in and immediately surrounding Boston (e.g., Quincy, Waltham, Somerville). Neighborhoods in Profile 7 were mostly in Western Massachusetts and the Cape Cod area (e.g. Franklin County, Provincetown). Neighborhoods in Profile 8 were disbursed in small pockets throughout the state (e.g., Peabody, Hyde Park, Tisbury, Adams). See Supplementary Table A.1 in the Supplementary Materials for more information about the census tracts population age and race-ethnicity distribution by NSDoH profile.
Figure 1: MBMM derived NSDoH profiles. Bars indicate the estimated posterior probability of high exposure to the NSDoH variable for a neighborhood, given assignment to an NSDoH profile. SDoH variables are ordered/colored by the NSDoH thematic domain. Abbreviations: high school(HS), EN(English), Non-Hispanic (NH), Supplemental Nutrition Assistance Program (SNAP).
Table 1: Distribution of selected neighborhood social determinants of health and ACS 2015-2019 table identification.
NSDoH ACS Variable (ACS Table) Median estimate for MA
% of owner-occupied housing (DP04) 66.15
% of renter-occupied housing (DP04) 33.85
% Households without a motor vehicle (DP04) 7.7
% Crowding in household (DP04)  [^1] Occupied housing units with 1.01 to 1.50 and 1.51 or more occupants per room/ All occupied housing units for the same calendar year[46]. 1.28
% Occupied housing units without complete plumbing (DP04) [^2] We binarized this variable differently because the median was 0. Therefore, census tracts for which there were no households lacking complete plumbing were categorized as ”0”, and the remaining tracts were categorized as ”1” to indicate lack of plumbing. 0
Estimate median household income in the past 12 months (inflation-adjusted; B19013) 82265
% Female Single-parent households with children younger than 18 (DP02) 3.9
%With Food Stamp/SNAP benefits in the past 12 months (DP03) 7.7
% Below 150% poverty level (C17002)  [^3] People living below 150% the poverty level/ All people for the same calendar year[46, 47]. 12.5
% Unemployed/unemployment rate (DP03) 2.9
% Employed population aged 16 years or older, working class (C24010) 53.9
% Population aged 25 years or older with less than 9 years of education (DP02) 2.5
% High School degree or higher (DP02) 93.5
% Bachelor’s degree or higher (DP02) 40.5
% Language other than English: Speak English less than ”very well”(DP02) 5.8
% Hispanics or Latinos (DP05) 5.95
% Non-Hispanic Black (DP05) 2.8
% Non-Hispanic Asian (DP05) 3.2
3.3 NSDoH Profiles and Receipt of Optimal Care for EC
Table B.1 shows sociodemographic and clinical information of the analytical sample of EC patients, who received optimal care or NCCN guidelines adherent treatment (82.3%). These patients were predominantly non-Hispanic White (87.2%), foreign-born (49.7%), aged 50-64 (46.1%), privately insured (46.8%), diagnosed at stage I(91.5%), and with grade 1 tumors (52.9%). Patients were initially treated at large medical facilities (75.7%), academic medical centers (38.7%), and diagnosed by physicians specializing in family/internal medicine (47%).
Table 2 details the distribution of patient demographics based on which NSDoH profile they resided at EC diagnosis. Race-ethnicity, birthplace, insurance status at diagnosis, and initial point of care facility characteristics were significantly associated with NSDoH profiles. For example, among patients residing in neighborhoods belonging to NSDoH profile pattern 2, 45.8% were foreign-born, 59.41% had Medicare or public insurance, and 39.4% were initially treated at academic medical centers.
Figure 2 summarizes the results of the regression analysis. Although not statistically significant, compared to patients who resided in neighborhoods belonging to the NSDoH Profile 1, patients in the NSDoH Profile 2 have lower odds [OR = 0.77, 95% Credible Interval (0.56,1.07)] of receiving optimal care after adjusting for year of diagnosis, age at diagnosis, insurance status at diagnosis, and initial type of care facility.
We also examined the relationship between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). The results, adjusted for year, age, and insurance status at diagnosis, are detailed in Supplementary Materials B. Compared to profile 1, patients who resided in NSDoH profiles 3 and 6 were significantly associated with higher odds of receiving care at an academic medical center, while those in NSDoH profile 5 had lower odds.
Figure 2: Adjusted odds of receiving optimal care by NSDoH profiles (Ref= NSDoH Profile 1) for patients diagnosed with EC from 2015-2019. Bars show 95% credible intervals.
4 Discussion
We used data from the ACS to derive NSDoH profiles using a fully Bayesian multivariate Bernoulli mixture model. To our knowledge, this is the first study to apply Bayesian mixture models to neighborhood-level data in this context. This approach allowed us to estimate the number of profiles directly from the data, eliminating the need for post-hoc testing. Our model identified eight (8) NSDoH profiles which we characterized using neighborhood SDoH variables belonging to four thematic domains: household conditions and resources, economic security, educational attainment, and social and community context.
We used these profiles to conduct a regression analysis where we examined the association of NSDoH profiles and optimal care for EC. Our findings, although not statistically significant, suggest disparities in EC care. Specifically, patients residing in the profile characterized as the most disadvantaged, based on the probability distribution of SDoH variables, had lower odds of receiving optimal care. These findings are comparable to Rodriguez et al. where they examined the association between neighborhood socioeconomic status (NSES) using the Yost Index in the Surveillance, Epidemiology, and End Results (SEER) database from 2006-2015 and adherence to the NCCN guidelines[12]. Rodriguez et al. was the first study, to our knowledge, to find a decreasing trend in adherence to NCCN guidelines for EC by decreasing level of NSES. Similarly to our findings, they observed that patients residing in the most disadvantaged neighborhoods with the lowest NSES had 0.73 times lower odds of receiving NCCN guideline-concordant care compared to those in the highest NSES group. In their study, 59.5% of patients received treatment adhering to NCCN guidelines, with the lowest adherence seen among Black, Latina, and American Indian/Alaska Native women (57.1%, 54.5%, and 52.7%, respectively). These proportions are lower than those in our study (82.3%) but comparable to a study using the Women’s Health Initiative (WHI), where they found 80% of patients received NCCN adherent treatment[48]. Our results focused specifically on Massachusetts, while their analysis covered the entire United States. Additionally, similar to the WHI study, our study population was predominantly Non-Hispanic White, which may have limited our ability to analyze disparities in optimal care among racial and ethnic minority groups. This is important, as differences in standard treatment regimens may exist for some groups due to variations in the histologic subtypes and stages of endometrial cancer they present with. [49] Lastly, beyond the limitations of retrospective observational studies, the MCR lacks information on comorbidities, which may also impact differences in treatment adherence and recommendations and may explain non-adherence rates among race-ethnic minority groups.
Rodriguez et al., along with other studies examining the impact of neighborhood socioeconomic status (SES) on cancer outcomes, have used the Yost Index from the SEER database [12, 11, 13, 50]. Initially developed in 2001 to examine breast cancer incidence in California, this measure of neighborhood disadvantage, although it uses a subset of our variables, provides different information and may not be reflective of healthcare access in the current socioeconomic context. This is because the Yost Index emphasizes monetary indicators and is derived via a one-component PCA. This method uses weights or the component loadings that explain the most variance within the data to create an uncorrelated linear combination of the SES indicator variables [51]. Although PCA is a powerful and widely used method to reduce the dimensionality of high-dimensional datasets, like in the case of neighborhood data, the method suffers from interpretability issues because it outputs a linear combination of the original indicators, and therefore we cannot generalize about the original variables [51]. In contrast, our approach using finite mixture models provides a holistic way to identify patterns within a dataset by modeling the data as a mixture of Bernoulli distributions. This approach provides interpretable clusters of neighborhoods making it easier to understand the heterogeneity of neighborhood data, as well as the defining characteristics of each profile. We used a Bayesian framework for mixture models because it provides the flexibility of estimating model parameters and the number of profiles simultaneously. Specifically, our approach treats the number of profiles like any other unknown parameter and puts a prior distribution on it, which provides a robust method for consistently estimating the number of profiles [52, 44, 53]. Utilizing a Bayesian framework to estimate our model allowed us to borrow information within and across other neighborhoods to generate these profiles and provide a more comprehensive view of how these factors interrelate with one another at the neighborhood level. Overall, using our NSDoH profiles to assess association with health outcomes, such as receipt of optimal care for EC, may give a more nuanced view of how SDoH work overlap and co-occur within communities rather than in isolation.
This model is still met with limitations. First, model-based clustering is reflective of the data we use. In this study, we used 5-year estimates from the ACS from 2015-2019 for MA. Thus, the current NSDoH profiles may vary across different ACS survey waves due to changes in the distribution of SDoH variables, as well as across different geographies. Our binarized thresholds were based on median values for the state of Massachusetts. A different state would yield different cutoffs and, ultimately, different profiles. Second, we made the decision to dichotomize the SDoH variables based on the skewed distributions found in the 2015-2019 cycles. This can sometimes result in a slight loss of information and reduced precision of the profile estimates since the full distribution of these variables was not considered. Future work should explore the development and implementation of a fully Bayesian multivariate beta mixture model that can flexibly accommodate bounded data, such as what we see in census-level data. Third, our analysis used aggregate data from the Census to construct NSDoH profiles. These variables are reflective of population-level characteristics rather than individual-level data. To avoid ecological fallacies, relationships observed at the aggregate level cannot be assumed to hold for individual EC patients residing within these neighborhoods. Lastly, consistent with other similar studies, our analysis is cross-sectional [23] and does not account for the changes in neighborhood demographics over time as the result of fair housing policies and gentrification. We focused our time frame on 2015-2019, which was based on the ACS 5-year estimates of data. Future studies should incorporate multiple survey cycles to assess how demographic shifts influence NSDoH profiles and, consequently, the access to and quality of healthcare and other outcomes for their residents.
5 Conclusion
Bayesian mixture models offer promising applications for neighborhood-level data. Our approach enabled us to characterize and assign NSDoH profile patterns to census tracts in Massachusetts. Geospatial mapping of NSDoH profiles demonstrated how we can leverage these tools to identify areas for targeted interventions. The profiles derived for the model were applied to examine the association with receipt of optimal care for endometrial cancer patients of certain NSDoH profiles. However, these profiles are translatable to a myriad of other outcomes and exploratory analyses.
Abbreviations
SDoH: Social determinants of health
NSDoH: Neighborhood social determinants of health
EC: Endometrial cancer
ACS: American Community Survey
MA: Massachusets
ADI: Area Deprivation Index
PCA: Principal Components Analysis
SES: Socioeconomic Status
FMM:Finite Mixture Models
LCA: Latent Class Analysis
LPA: Latent Profile Analysis
MBMM: Multivariate Bernoulli Mixture Model
SNAP: Supplemental Nutrition Assistance Program
GED: General Education Development
MC3: Metropolis-coupled Markov chain Monte Carlo algorithm
MCR: Massachusetts Cancer Registry
MDPH: Massachusetts Department of Public Health
NCCN: National Comprehensive Cancer Network
Declarations
Funding
This project was supported by the Harvard Data Science Initiative Bias2 program. Dr. Stephenson is supported by NHLBI 1K01HL166442-01. Carmen Rodriguez Cabrera is supported by NIH T32CA009337. The funders had no role in the conceptualization, design, data analysis, decision to publish, or preparation of the manuscript.
Competing interests
The authors declare that they have no competing interests.
Ethics approval and consent to participate
The institutional review board at the Massachusetts Department of Public Health (MDPH) approved the ethical clearance for the MCR-MDPH data in partnership with Harvard University School of Public Health.
Consent for publication
Not applicable.
Availability of data and materials
The R code used for this manuscript has been made publicly available in GitHub, https://github.com/cbrodriguez01/ecbayesbinmix. This also includes all the American Community Survey data we used for multivariate Bernoulli mixture model, which is also publicly available at https://www.census.gov/programs-surveys/acs/data.html. The data that support the findings from the regression analysis in this study are available from the Massachusetts Cancer Registry-Massachusetts Department of Public Health (MCR-MDPH), but restrictions apply to the availability MCR data, which were used under license for the current study, and so are not publicly available. Data are, however, available from the authors upon reasonable request and with permission of MCR-MDPH.
Author information
Authors and Affiliations
Department of Biostatistics, Harvard T.H. Chan School of Public Health, Boston, Massachusetts
Carmen B. Rodriguez, Stephanie M. Wu, and Briana J.K. Stephenson
Division of Gynecologic Oncology, Department of Obstetrics and Gynecology and Reproductive Biology, Brigham and Women’s Hospital, Boston, Massachusetts
Harvard Medical School, Boston, Massachusetts
Stephanie Alimena
Dana-Farber Cancer Institute, Boston, Massachusetts
Carmen B. Rodriguez and Stephanie Alimena
Department of Health Policy and Management, Harvard T.H. Chan School of Public Health, Boston, Massachusetts
Alecia J. McGregor
Contributions
CBR and BJKS conceptualized and designed the study. CBR performed all the data wrangling, analysis, interpretation of results, and drafting of the manuscript. SMW, SA, and AJM assisted in the interpretation of results. BJKS provided input and supervision over all data analysis and interpretation, as well as provided critical review and revisions to the manuscript. All authors read, contributed, and approved the final manuscript.
Corresponding author
Correspondence to Briana J.K. Stephenson.
Acknowledgments
We are grateful to everyone who contributed ideas or feedback, whether through brief conversations or informal discussions. We would also like to thank Dr. Giovanni Parmigiani for his valuable insight.
Table 2: Sociodemographic and Clinical Characteristics of Endometrial Carcinoma Cases Between 2015 – 2017 in the Massachusetts Cancer Registry (MCR) by NSDoH profiles (n=2412).
Neighborhood SDoH Profiles
n(%) Profile 1 Profile 2 Profile 3 Profile 4 Profile 5 Profile 6 Profile 7 Profile 8
n=817 n= 426 n=236 n=285 n=196 n= 130 n=179 n=143
Optimal care status
Not optimal 137 (16.8) 89 (20.9) 42 (17.8) 51 (17.9) 32 (16.3) 18 (13.8) 33 (18.4) 25 (17.5)
Optimal 680 (83.2) 337 (79.1) 194 (82.2) 234 (82.1) 164 (83.7) 112 (86.2) 146 (81.6) 118 (82.5)
Race-Ethnicity
Non-Hispanic White 776 (95.0) 275 (64.6) 213 (90.3) 271 (95.1) 170 (86.7) 94 (72.3) 169 (94.4) 135 (94.4)
Non-Hispanic Black 10 ( 1.2) 52 (12.2) 4 ( 1.7) 0 ( 0.0) 13 ( 6.6) 13 (10.0) 4 ( 2.2) 3 ( 2.1)
Hispanic 22 ( 2.7) 16 ( 3.8) 7 ( 3.0) 7 ( 2.5) 1 ( 0.5) 14 (10.8) 3 ( 1.7) 2 ( 1.4)
Other 7 ( 0.9) 78 (18.3) 11 ( 4.7) 6 ( 2.1) 9 ( 4.6) 6 ( 4.6) 1 ( 0.6) 1 ( 0.7)
Birthplace
US-born 391 (47.9) 136 (31.9) 123 (52.1) 112 (39.3) 60 (30.6) 42 (32.3) 66 (36.9) 38 (26.6)
Foreign-born 373 (45.7) 195 (45.8) 98 (41.5) 158 (55.4) 112 (57.1) 59 (45.4) 106 (59.2) 100 (69.9)
Unknown 53 ( 6.5) 95 (22.3) 15 ( 6.4) 15 ( 5.3) 24 (12.2) 29 (22.3) 7 ( 3.9) 5 ( 3.5)
Year of diagnosis (y)
2015 270 (33.0) 156 (36.6) 90 (38.1) 87 (30.5) 58 (29.6) 47 (36.2) 67 (37.4) 46 (32.2)
2016 296 (36.2) 153 (35.9) 75 (31.8) 111 (38.9) 72 (36.7) 45 (34.6) 56 (31.3) 47 (32.9)
2017 251 (30.7) 117 (27.5) 71 (30.1) 87 (30.5) 66 (33.7) 38 (29.2) 56 (31.3) 50 (35.0)
Age at diagnosis (y)
Younger than 50 58 ( 7.1) 52 (12.2) 29 (12.3) 21 ( 7.4) 19 ( 9.7) 16 (12.3) 10 ( 5.6) 9 ( 6.3)
50-64 366 (44.8) 198 (46.5) 111 (47.0) 132 (46.3) 90 (45.9) 58 (44.6) 84 (46.9) 72 (50.3)
65 or older 393 (48.1) 176 (41.3) 96 (40.7) 132 (46.3) 87 (44.4) 56 (43.1) 85 (47.5) 62 (43.4)
Insurance status at diagnosis
Private 405 (49.6) 130 (30.5) 118 (50.0) 136 (47.7) 84 (42.9) 62 (47.7) 70 (39.1) 59 (41.3)
Medicare 341 (41.7) 175 (41.1) 83 (35.2) 123 (43.2) 76 (38.8) 49 (37.7) 81 (45.3) 58 (40.6)
Public/Government 35 ( 4.3) 78 (18.3) 19 ( 8.1) 14 ( 4.9) 16 ( 8.2) 8 ( 6.2) 16 ( 8.9) 13 ( 9.1)
Other 22 ( 2.7) 39 ( 9.2) 13 ( 5.5) 9 ( 3.2) 18 ( 9.2) 8 ( 6.2) 10 ( 5.6) 12 ( 8.4)
Not insured 14 ( 1.7) 4 ( 0.9) 3 ( 1.3) 3 ( 1.1) 2 ( 1.0) 3 ( 2.3) 2 ( 1.1) 1 ( 0.7)
Type of surgery received
None 24 ( 2.9) 18 ( 4.2) 9 ( 3.8) 13 ( 4.6) 8 ( 4.1) 6 ( 4.6) 7 ( 3.9) 6 ( 4.2)
Resection 792 (96.9) 403 (94.6) 227 (96.2) 269 (94.4) 188 (95.9) 124 (95.4) 169 (94.4) 137 (95.8)
Other Surgery/Unknown 1 ( 0.1) 5 ( 1.2) 0 ( 0.0) 3 ( 1.1) 0 ( 0.0) 0 ( 0.0) 3 ( 1.7) 0 ( 0.0)
Type of radiation administered
No radiation treatment 591 (72.3) 310 (72.8) 180 (76.3) 216 (75.8) 142 (72.4) 104 (80.0) 133 (74.3) 113 (79.0)
External beam radiation therapy (EBRT) 5 ( 0.6) 5 ( 1.2) 1 ( 0.4) 2 ( 0.7) 0 ( 0.0) 4 ( 3.1) 0 ( 0.0) 1 ( 0.7)
Brachytherapy 149 (18.2) 66 (15.5) 33 (14.0) 43 (15.1) 31 (15.8) 12 ( 9.2) 32 (17.9) 17 (11.9)
Other 72 ( 8.8) 45 (10.6) 22 ( 9.3) 24 ( 8.4) 23 (11.7) 10 ( 7.7) 14 ( 7.8) 12 ( 8.4)
Chemotherapy status
No 699 (85.6) 366 (85.9) 206 (87.3) 261 (91.6) 172 (87.8) 109 (83.8) 165 (92.2) 123 (86.0)
Yes 118 (14.4) 60 (14.1) 30 (12.7) 24 ( 8.4) 24 (12.2) 21 (16.2) 14 ( 7.8) 20 (14.0)
Stage at diagnosis
Stage I 687 (84.1) 345 (81.0) 186 (78.8) 246 (86.3) 164 (83.7) 105 (80.8) 151 (84.4) 119 (83.2)
Stage II 78 ( 9.5) 44 (10.3) 30 (12.7) 27 ( 9.5) 15 ( 7.7) 13 (10.0) 15 ( 8.4) 14 ( 9.8)
Stage III 27 ( 3.3) 12 ( 2.8) 11 ( 4.7) 5 ( 1.8) 10 ( 5.1) 5 ( 3.8) 2 ( 1.1) 4 ( 2.8)
Stage IV 25 ( 3.1) 25 ( 5.9) 9 ( 3.8) 7 ( 2.5) 7 ( 3.6) 7 ( 5.4) 11 ( 6.1) 6 ( 4.2)
Grade at diagnosis
Grade 1 371 (45.4) 217 (50.9) 103 (43.6) 149 (52.3) 94 (48.0) 63 (48.5) 91 (50.8) 78 (54.5)
Grade 2 284 (34.8) 122 (28.6) 89 (37.7) 97 (34.0) 71 (36.2) 39 (30.0) 57 (31.8) 44 (30.8)
Grade 3 162 (19.8) 87 (20.4) 44 (18.6) 39 (13.7) 31 (15.8) 28 (21.5) 31 (17.3) 21 (14.7)
Initial point of care facility type
Academic Medical Centers 318 (38.9) 168 (39.4) 113 (47.9) 104 (36.5) 53 (27.0) 71 (54.6) 68 (38.0) 48 (33.6)
Community 336 (41.1) 124 (29.1) 68 (28.8) 106 (37.2) 69 (35.2) 27 (20.8) 44 (24.6) 48 (33.6)
Specialty 16 ( 2.0) 5 ( 1.2) 1 ( 0.4) 2 ( 0.7) 2 ( 1.0) 3 ( 2.3) 1 ( 0.6) 1 ( 0.7)
Teaching 132 (16.2) 110 (25.8) 54 (22.9) 61 (21.4) 66 (33.7) 29 (22.3) 60 (33.5) 44 (30.8)
Initial point of care facility size
Small (<100) 21 ( 2.6) 7 ( 1.6) 3 ( 1.3) 11 ( 3.9) 2 ( 1.0) 1 ( 0.8) 13 ( 7.3) 3 ( 2.1)
Medium (100-299) 213 (26.1) 71 (16.7) 47 (19.9) 72 (25.3) 32 (16.3) 22 (16.9) 36 (20.1) 33 (23.1)
Large (300+) 583 (71.4) 348 (81.7) 186 (78.8) 202 (70.9) 162 (82.7) 107 (82.3) 130 (72.6) 107 (74.8)
Initial point of care facility , doctor specialty
Family/Internal medicine 427 (52.3) 192 (45.1) 138 (58.5) 105 (36.8) 81 (41.3) 80 (61.5) 43 (24.0) 52 (36.4)
Hematology 13 ( 1.6) 3 ( 0.7) 3 ( 1.3) 3 ( 1.1) 0 ( 0.0) 0 ( 0.0) 1 ( 0.6) 1 ( 0.7)
Gynecology and obstetrics 155 (19.0) 83 (19.5) 39 (16.5) 71 (24.9) 48 (24.5) 13 (10.0) 46 (25.7) 27 (18.9)
Oncology 105 (12.9) 63 (14.8) 28 (11.9) 39 (13.7) 31 (15.8) 18 (13.8) 35 (19.6) 29 (20.3)
Radiology 25 ( 3.1) 10 ( 2.3) 4 ( 1.7) 5 ( 1.8) 5 ( 2.6) 3 ( 2.3) 7 ( 3.9) 5 ( 3.5)
Other specialty 15 ( 1.8) 7 ( 1.6) 6 ( 2.5) 8 ( 2.8) 4 ( 2.0) 4 ( 3.1) 5 ( 2.8) 3 ( 2.1)
Unknown 77 ( 9.4) 68 (16.0) 18 ( 7.6) 54 (18.9) 27 (13.8) 12 ( 9.2) 42 (23.5) 26 (18.2)
Supporting Materials
Appendix A Multivariate Bernoulli Mixture Model
Figure A.1 shows the pairwise correlations of the selected NSDoH variables. Given that we used a small geographic unit (i.e., census tracts), some variables are highly correlated. In the case of the MBMM, the variance-covariance matrix of a mixture component of independent Bernoulli distributions is not diagonal and does not assume independence between variables within clusters; therefore, this model can accommodate neighborhood-level data.
Figure A.1: Pairwise Pearson’s Correlations of NSDoH variables from ACS 2015-2019.
We assumed that the observed data comes from a mixture of $K$ independent Bernoulli distributions. Here, $K$ is the number of mixture components (i.e., clusters) [^1] In the manuscript text we refer to clusters as profiles. included in the model, while the true number of clusters in the data is $K_{0}<K$ and is determined by model estimation. That is, we assume that for a binary data matrix $\mathbf{X}=\{\mathbf{x_{1}},...\mathbf{x_{n}}\}$, each $\mathbf{x_{i}}=\{x_{i,1},...,x_{i,p}\}$ is such that $p(x_{i})=\sum_{k=1}^{K}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}$ as described in the methods section 2.2, where $\boldsymbol{\pi}=\{\pi_{1},...,\pi_{K}\}$ is the probability vector for the cluster assignments (i.e., the probability that a census tract belongs to cluster $k\in\{1,...,K\}$) and $\sum_{k=1}^{K}\pi_{k}=1$, and the probability matrix $\boldsymbol{\theta}=\{\theta_{j|k}\}^{p\times K}$ represents the probability of a high level of exposure to NSDoH variable $j$ for a census tract given assignment to cluster $k$. Therefore, the observed likelihood for the specified model is
$\mathcal{L}(\boldsymbol{\pi},\boldsymbol{\theta}|\mathbf{X})=\prod_{i=1}^{n}\sum_{k=1}^{K}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}$ (3)
We augment the data by introducing a latent allocation variable $\boldsymbol{z_{i}}$, such that $z_{i}=k$ means that census tract $i$ has been generated from the $k$-th cluster, and thus $P(z_{i}=k)=\pi_{k}$. For inference, we consider the complete data $\{x_{i},z_{i}\}$ likelihood for the MBMM:
$\mathcal{L}^{c}(\boldsymbol{\pi},\boldsymbol{\theta}|\mathbf{X},\mathbf{Z})=\prod_{i=1}^{n}\prod_{k=1}^{K}\biggl{\{}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}\biggl{\}}^{\mathbb{I}(z_{i}=k)}$ (4)
The Bayesian framework for computation relies on three key data components: prior information about the model parameters, observed data likelihood, and posterior information. The prior information is defined as the first estimates of the distribution of the parameters before incorporating data through the Markov chain Monte Carlo (MCMC) sampler. As the sampler advances, both prior and observed data synergistically contribute to formulating insights about our target posterior distribution and updating information about the parameters using Bayes’ Theorem.
Estimation of the parameters for the MBMM was performed using a Bayesian sampler described and implemented by Panagiotis Papastamoulis and Magnus Rattray as the R package BayesbinMix [52]. The main function in the package is coupleMetropolis(), which embeds an allocation sampler [54] with an unknown number of mixture components (i.e., a way to estimate the optimal number of components simultaneously) in a Metropolis-coupled Markov chain Monte Carlo ($MC^{3}$) algorithm. The $MC^{3}$ strategy is adopted to improve MCMC sampling by considering heated versions of the original target distribution. The function coupleMetropolis() takes as input the binarized observed data matrix $\mathbf{X}$, an upper bound on the number of clusters, defined in the function as $K_{max}$, prior information for all model parameters, and other inputs for computational efficiency. It then outputs the estimated posterior distribution of the model parameters $(\boldsymbol{\pi},\boldsymbol{\theta})$ and the most probable number of NSDoH profiles defined as $(K_{map})$.For detailed information about this approach, please see their article BayesBinMix: an R Package for Model-Based Clustering of Multivariate Binary Data [52].
We assume no prior knowledge for the initialization of the parameters. Consequently, all parameters are initialized with non-informative priors detailed below. The true number of NSDoH profiles is unknown. Using the upper bound $K_{max}$( also referred to as $K$ in our model description above), we impose an overfitted finite mixture model[44]. We fit the model with a large upper bound on the number of clusters, $K=50$, coupled with a prior on the number of clusters $K<K_{max}$, where $K$ is treated by the model as another unknown parameter, and allow a data-driven approach to estimating $K$[44, 24]. We impose the following priors on model parameters:
$K|K_{max}\sim\text{Poisson}(\lambda=1)\text{ truncated on the set}\{1,...,K_{max}\}$
For the other model parameters, we assume the following priors:
$\boldsymbol{\pi}\mid K\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{K_{m}ax}),\quad\text{where }\gamma_{k}=1\;\forall\;k.$
$\theta_{j|k}|K\sim Beta(\alpha,\beta)\quad\text{where }\alpha=1=\beta\;\forall\;j,k$
Additionally, we conducted sensitivity analyses using other priors allowed in the software package. Namely, we used $K\sim Uniform(1,K_{max})$, and $\boldsymbol{\pi}|K\sim\text{Dirichlet}(\frac{1}{K_{max}},...,\frac{1}{K_{max}})$, but we found no significant differences, in fact, the model behaved better with the priors specified above. The following full conditional distributions were used to update the model parameters:
$\boldsymbol{\pi}|K,\mathbf{Z}\sim\text{Dirichlet}\bigg{(}\gamma_{1}+\sum_{i=1}^{n}\mathbb{I}(z_{i}=1),...,\gamma_{K}+\sum_{i=1}^{n}\mathbb{I}(z_{i}=K)\bigg{)}$
$\theta_{j|k}|K,\mathbf{X},\mathbf{Z}\sim\text{Beta}\bigg{(}\alpha+\sum_{i=1}^{n}\mathbb{I}(z_{i}=k)x_{i,j},\beta+\sum_{i=1}^{n}\mathbb{I}(z_{i}=k)-\sum_{i=1}^{n}\mathbb{I}(z_{i}=k)x_{i,j}\bigg{)}$
$P(z_{i}=k|K,\mathbf{x}_{i},\boldsymbol{\pi},\boldsymbol{\theta})\propto\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}$
Given that they are smaller geographic units, some census tracts have missing information on some of the NSDoH variables (ranging from 14-24 census tracts with missing across all variables). This model can handle missing data by imputing these values using the parameter’s posterior mean estimates. The $MC^{3}$ for posterior computation was run for 15,000 iterations, with thinning every 10 iterations (i.e., retaining every 10th sample) and the first 5000 iterations removed as part of posterior samples post-processing. For the MCMC chain heating parameter denoted as $h_{m},m=2,...,M$, where $M$ is the total number of parallel chains, we used incremental heating where the heat of the $m$’th chain is $h_{m}=1/[1+\Delta T\times(m-1)]$, and we tuned the parameter $\Delta T$ such that swaps between chains were accepted 20%- 60% of time [55]. Through this process, we found that four (4) heated chains for the $MC^{3}$ algorithm produced good mixing, and for this type of data, smaller $\Delta T=0.025$ works better. Posterior mean estimates were calculated from the remaining 1000 iterations collected from the sampler’s output. The generated MCMC samples were postprocessed using the Equivalence Classes Representatives (ECR) algorithm to overcome label-switching identifiability issues inherent in Bayesian mixture models[56]. The most probable number of clusters ($K_{map}$) given the data is inferred, and accordingly, the NSDoH profile assignment probabilities for each census tract are estimated after reordering with the ECR algorithm given $K_{map}$. Figure A.2 shows the distribution of the assignment probabilities for all NSDoH profiles given $K_{map}$ from the model.
The model yielded nine (9) NSDoH profiles. However, we applied additional post-processing steps to reclassify census tracts that were part of smaller NSDoH profiles, defined as profiles containing $<$ 5% of the total census tracts. This step was taken to produce more meaningful and interpretable clusters. Reclassification was based on the second-highest posterior probability, as it was not significantly lower than the highest probability. Additionally, as shown in Figure A.2, some cluster assignment probabilities were low, which may suggest that there is unobserved heterogeneity (e.g., due to unobserved NSDoH variables), which the model did not account for. The final clusters were qualitatively described based on thematic domains to create NSDoH profiles as discussed in Section 3.1.
Figure A.2: Distribution of assignment probabilities for each NSDoH profile.
Table A.1 shows the distribution of age and race/ethnicity within the census tracts used to further characterize NSDoH profiles in Section 3.2.
Table A.1: Distribution of population characteristics within census tracts in a given NSDoH Profile.
Neighborhood SDoH Profiles
Profile 1 Profile 2 Profile 3 Profile 4 Profile 5 Profile 6 Profile 7 Profile 8
Mean (SD) n = 401 n = 378 n= 185 n = 142 n= 109 n = 100 n = 87 n = 76
Census tract population 5090 (1860) 4220 (1690) 4300 (1730) 5000 (1820) 4910 (1870) 4700 (1650) 4140 (1560) 4550 (1990)
Median Age 44.7 (6.24) 35.3 (5.23) 35.3 (6.76) 46.2 (4.72) 42.1 (5.07) 34.5 (5.13) 46.6 (8.33) 41.6 (7.02)
Ages 20-24 5.36 (5.26) 7.96 (4.29) 10.4 (9.39) 5.47 (1.67) 7.12 (6.28) 9.92 (8.06) 6.45 (6.35) 7.32 (8.07)
Ages 25-34 8.94 (3.57) 16.4 (4.37) 24.4 (12.3) 10.7 (2.51) 12.4 (3.57) 25.4 (7.94) 10.6 (3.97) 14.2 (4.55)
Ages 35-44 11.6 (2.89) 13.1 (3.25) 12.6 (3.75) 11.2 (2.33) 12.3 (3.19) 13.1 (3.87) 10.3 (3.03) 12.1 (3.72)
Age 45-54 15.5 (3.21) 12.6 (3.10) 10.9 (4.45) 15.1 (2.52) 14.2 (3.23) 10.7 (3.56) 13.2 (3.45) 12.9 (3.48)
Age 55-59 8.07 (2.05) 6.09 (1.84) 5.30 (2.52) 8.52 (1.85) 7.91 (1.98) 5.60 (2.39) 7.84 (2.49) 7.65 (3.65)
Ages 60-64 7.21 (2.31) 5.44 (2.06) 4.86 (2.29) 8.16 (1.79) 7.21 (2.20) 5.03 (2.12) 8.23 (2.79) 7.05 (2.60)
Ages 65-74 11.0 (4.80) 7.17 (2.61) 7.51 (3.40) 11.8 (3.34) 9.50 (2.55) 7.21 (2.97) 13.2 (5.37) 9.83 (3.66)
Ages 75-84 5.43 (2.32) 3.60 (1.93) 3.78 (2.23) 5.48 (2.61) 4.87 (2.02) 3.73 (1.83) 6.47 (3.08) 5.02 (2.64)
Age 85 or older 2.60 (1.81) 1.78 (1.31) 1.99 (1.91) 2.43 (1.81) 2.56 (1.72) 2.01 (1.75) 2.98 (1.74) 2.69 (2.44)
Non-Hispanic White 85.9 (8.81) 41.9 (24.8) 72.9 (12.0) 92.2 (4.52) 71.8 (18.8) 57.4 (12.9) 86.7 (9.43) 81.6 (15.1)
Non-Hispanic Black 2.18 (3.54) 16.1 (18.9) 4.36 (4.09) 1.29 (1.40) 9.19 (12.0) 10.1 (8.27) 3.35 (4.20) 4.73 (7.07)
Non-Hispanic Asian 6.26 (6.32) 5.79 (9.00) 12.1 (8.87) 1.41 (1.81) 3.81 (4.16) 15.0 (10.1) 2.51 (2.53) 1.65 (2.44)
Hispanic or Latino 3.40 (2.48) 31.7 (22.4) 7.17 (4.00) 3.08 (2.91) 11.6 (9.70) 13.9 (10.5) 4.71 (4.40) 8.91 (12.4)
Appendix B Regression Analysis Additional
The main outcome is optimal care, defined as adherence to National Comprehensive Cancer Network (NCCN) guidelines. These guidelines recommend a combination of therapies depending on the stage and grade of the tumor (Figure B.1). We used year-specific guidelines spanning the study period due to gradual changes in NCCN treatment guidelines. We determined optimal care by comparing the treatment received and the treatment recommended by NCCN. For example, we know that surgery is the first course of treatment, and thus if a woman received surgery alone or surgery and other additional therapies, as shown in Figure B.1, based on the stage and grade of their tumor, then she is classified as having received optimal care. We combined all the corresponding data to create a binary outcome variable, where:
$Y_{i}=\begin{cases}1&\text{Received optimal care: patient received treatment following NCCN guidelines}\\ 0&\text{Did not receive optimal care}\end{cases}$
Figure B.1: National Comprehensive Cancer Network guidelines example from 2020.
Patient’s sociodemographic characteristics included age at diagnosis (younger than 50 years old, 50-64 years, or 65 and older), health insurance status at diagnosis (private, Medicare, public/government, uninsured/other), nativity ( U.S vs. foreign-born), and race/ethnicity (Hispanic, Non-Hispanic White, Non-Hispanic Black, Other race/ethnicity). Clinical characteristics included summary stage (distant, localized, regional), FIGO-stage (I,II,III,IV) using translations from the Summary Stage 2018 Coding Manual, tumor grade (1,2,3), and year of diagnosis. Treatment variables included type of surgery( none, resection, tumor destruction, other/unknown), type of radiation (brachytherapy, external beam (EBRT), other), and chemotherapy status (no, yes), and dates for each corresponding procedure. Additionally, we included information on the type (academic medical centers, community, specialty, and teaching) and size of the facility at the initial point of cancer management.
Table B.1 shows the distribution of patient characteristics by optimal care status defined as NCCN treatment adherence.
Table B.1: Sociodemographic and Clinical Characteristics of Massachusetts Cancer Registry Endometrial Cancer Cases from 2015-2017 by Optimal Care Status (n=2412).
Received
optimal care
(n=1985)
Did not receive
optimal care
(n= 427)
Race-Ethnicity
Non-Hispanic White 1739 (87.6) 364 (85.2)
Non-Hispanic Black 75 ( 3.8) 24 ( 5.6)
Hispanic 58 ( 2.9) 14 ( 3.3)
Other 99 ( 5.0) 20 ( 4.7)
Birthplace
US-born 796 (40.1) 172 (40.3)
Foreign-born 986 (49.7) 215 (50.4)
Unknown 203 (10.2) 40 ( 9.4)
Year of diagnosis (y)
2015 667 (33.6) 154 (36.1)
2016 723 (36.4) 132 (30.9)
2017 595 (30.0) 141 (33.0)
Age at diagnosis (y)
Younger than 50 181 ( 9.1) 33 ( 7.7)
50-64 954 (48.1) 157 (36.8)
65 or older 850 (42.8) 237 (55.5)
Insurance Status at Diagnosis
Private 928 (46.8) 136 (31.9)
Medicare 769 (38.7) 217 (50.8)
Public/Government 152 ( 7.7) 47 (11.0)
Other 110 ( 5.5) 21 ( 4.9)
Not insured 26 ( 1.3) 6 ( 1.4)
Type of surgery received
None 0 ( 0.0) 91 (21.3)
Resection 1983 (99.9) 326 (76.3)
Other/Unknown 2 ( 0.1) 10 ( 2.3)
Type of radiation administered
No radiation treatment 1558 (78.5) 231 (54.1)
External beam (EBRT) 14 ( 0.7) 4 ( 0.9)
Brachytherapy 367 (18.5) 16 ( 3.7)
Other 46 ( 2.3) 176 (41.2)
Chemotherapy status
No 1828 (92.1) 273 (63.9)
Yes 157 ( 7.9) 154 (36.1)
Stage at diagnosis
Stage I 1816 (91.5) 187 (43.8)
Stage II 54 ( 2.7) 182 (42.6)
Stage III 62 ( 3.1) 14 ( 3.3)
Stage IV 53 ( 2.7) 44 (10.3)
Grade at diagnosis
Grade 1 1050 (52.9) 116 (27.2)
Grade 2 670 (33.8) 133 (31.1)
Grade 3 265 (13.4) 178 (41.7)
Initial point of care facility type
Academic Medical Centers 768 (38.7) 175 (41.0)
Community 693 (34.9) 129 (30.2)
Specialty 15 ( 0.8) 16 ( 3.7)
Teaching 469 (23.6) 87 (20.4)
Initial point of care facility size
[…]
Medium (100-299) 450 (22.7) 76 (17.8)
Large (300+) 1496 (75.4) 329 (77.0)
Initial point of care facility , doctor specialty
Family/Internal medicine 933 (47.0) 185 (43.3)
Hematology 12 ( 0.6) 12 ( 2.8)
Gynecology & obstetrics 407 (20.5) 75 (17.6)
Oncology 305 (15.4) 43 (10.1)
Radiology 45 ( 2.3) 19 ( 4.4)
Other specialty 44 ( 2.2) 8 ( 1.9)
Unknown 239 (12.0) 85 (19.9)
We also explored the association between NSDoH profiles and the type of facility where patients were diagnosed or received treatment. Due to data limitations, we could not distinguish between these scenarios. For this sub-analysis, we dichotomized the facility type as academic medical center versus other types. Similarly to the optimal care analysis, we conducted a logistic regression adjusted for age, year, and insurance status at diagnosis, with results presented in Table B.2. Even after adjustment, NSDoH profiles 3,5, and 6 were significantly associated with receiving care at an academic medical center (vs. other). NSDoH profiles 3 and 6 were linked to higher odds of receiving care at academic centers, while profile 5 was associated with lower odds. The interactive map in 3.2 shows that profile 3 neighborhoods are often located near universities, which are in proximity to academic medical centers. In contrast, patients residing in neighborhoods in NSDoH profile 5 received care at either a community or teaching hospital (Table 2), and from observing the interactive map in 3.2 for NSDoH profile 5, most neighborhoods may be located farther from academic medical centers, making it less convenient for residents to access these facilities. Transportation challenges, despite lower probabilities of ”no vehicle in the household” in this profile, may still be a barrier, particularly for working-class populations (for which there is a high probability of 79%) with limited flexibility in work schedules.
Table B.2: Univariate and Multivariate Logistic Regression of Receiving Care at an Academic Medical Center Among Endometrial Cancer Cases Between 2015 and 2017 in the MCR (n=2412).
Model $1$Unadjusted model. Model $2$Adjusted for age, year, and insurance status at diagnosis.
OR 95% Credible Interval OR 95% Credible Interval
Neighborhood SDoH Profile
Profile 1 Referent Referent
Profile 2 1.018 (0.792,1.311) 1.096 (0.856, 1.411)
Profile 3 1.439 (1.068,1.94) 1.452 (1.081, 1.954)
Profile 4 0.9 (0.683,1.178) 0.902 (0.679, 1.186)
Profile 5 0.577 (0.403,0.81) 0.588 (0.416, 0.823)
Profile 6 1.886 (1.325,2.671) 1.907 (1.299, 2.799)
Profile 7 0.959 (0.688,1.341) 1.001 (0.712, 1.413)
Profile 8 0.786 (0.545,1.131) 0.821 (0.561, 1.190)